# Direct Imitation Learning-based Visual Servoing using the Large Projection Formulation

### v1

@@ Abstract
Today robots must be safe, versatile, and user-friendly to operate in unstructured and human-populated environments. Dynamical system-based imitation learning enables robots to perform complex tasks stably and without explicit programming, greatly simplifying their real-world deployment. To exploit the full potential of these systems it is crucial to implement closed loops that use visual feedback. Vision permits to cope with environmental changes, but is complex to handle due to the high dimension of the image space. This study introduces a dynamical system-based imitation learning for direct visual servoing. It leverages off-the-shelf deep learning-based perception backbones to extract robust features from the raw input image, and an imitation learning strategy to execute sophisticated robot motions. The learning blocks are integrated using the large projection task priority formulation. As demonstrated through extensive experimental analysis, the proposed method realizes complex tasks with a robotic manipulator.
@@
Figure 1: Our work combines off-the-shelf deep learning strategies to detect objects in the clutter, and imitation learning to realize complicated trajectories, e.g., dropping a cube into a cup on an untidy table. The large projection formulation combines the two ML components and ensures convergence to a given target.
@@ 1 Introduction
Modern robots must be accessible to everyone, as they are rapidly spreading in everyday life environments, from industries [1] to hotels [2] and hospitals [3]. It is expected that a growing number of inexperienced end-users, like children, patients, or elderly people, ask for robots with easy-to-use, friendly, and modular interfaces, endowed with adaptive skills. To meet these requirements, recent advancements in robotics demonstrate the great potential of Machine Learning (ML).
@@ 2 Related work
One possibility consists of coupling perception and control together in end-to-end learning fashion [29, 30, 31]. However, end-to-end methods cannot guarantee stability properties and robustness to disturbances. Coarse-to-fine imitation learning [32] combines closed- and open-loop execution to perform complex manipulation tasks. One way to ensure stability is to maintain the formal structure of the visual controller, e.g., preserving the VS formalism. In the context of our work, Direct (DVS) is particularly interesting because it implements VS using direct image measurement, avoiding explicit feature extraction. Examples are VS schemes that use photometric moments [33], pixel luminance [34], histograms [35], or Gaussian mixtures [36] as control feedback. However, in these approaches, the potential of DL is not fully exploited. In [37], an artificial Neural Network (NN) is trained using the knowledge of VS to produce geometrically interpretable features. In [38], an autoencoder is learned to reduce the dimensionality of the image space, and an interaction matrix is directly computed from the network using auto differentiation and utilized in a VS law. Liu et al. [39] simplify the YOLO backbone to speed up the object detection, while Luo et al. [40] propose a top-down feature detection network, and both use the predicted features in a VS scheme. These approaches perform well with VS tasks from images without performing classical feature extraction but do not execute complex movements.
This paper proposes an approach to solve the perception problem and the generation of complex trajectories simultaneously, in the context of vision-based controllers. The proposed framework overcomes the limitations of the literature by proposing a Direct Imitation Learning-based Visual Servoing (D-ILVS) strategy that integrates off-the-shelf DL-based perception backbones with IL in a control theoretic framework.
@@ 3 Background
Image-based VS [10] is a more than twenty-year-old established technique to regulate a camera to a desired pose through visual information. The most basic law computes 6D velocity commands as $\bm{v}=-\lambda\hat{\bm{L}}^{+}\bm{e}$ [10], zeroing an error $\bm{e}\in\mathbb{R}^{f}$ defined on the image; $\lambda$ is a scalar gain; $\hat{\bm{L}}^{+}\in\mathbb{R}^{f\times 6}$ is the pseudo-inverse of the so-called interaction matrix, relating the camera velocity to the time derivative of the visual feedback; the hat over the matrix indicates the approximation due to unknown $3$D parameter.
Figure 2: State-of-the-art DL-based systems like YOLO can be used to detect the features of an object of interest on raw image robustly. Examples of features are the vertices (denoted with C1, C2, C3, and C4) of the bounding box detected around the image of the object. However, such detection systems fail to capture the correct object orientation: in the three snapshots, YOLO provides very similar feature values that correspond to very different relative camera-object orientations.
@@
$\bm{v}_{e}=-\lambda\hat{\bm{L}}^{+}\bm{e}+\bm{P}\bm{\sigma}$ (1)
where $\bm{P}=\bm{I}_{6}-\hat{\bm{L}}^{+}\hat{\bm{L}}$, and $\bm{\sigma}\in\mathbb{R}^{6}$ is the desired velocity realizing the secondary task. The main limitation of (1) is that, in normal working conditions, the dimension of the feedback is always greater or equal to the dimension of the task (i.e., $f\geq 6$[10]. Under these circumstances, there is not much room in the null space of the primary task to execute any other secondary task. Therefore, it has been proposed to use the norm of the error $\eta=\|\bm{e}\|$ as the primary task [42], and to realize a prioritized control scheme in the following form
@@
where the matrices of interest can be retrieved in closed form. Note that the control law (2) drives the system towards the original desired behavior, as $\bm{e}\to\bm{0}$ for $\eta\to 0$. The operator $\bm{P}_{\eta}$ is called large projector as the law (2) ensures enough room in the null space to achieve secondary tasks. In practice, regulating the scalar norm instead of the vector error releases degrees of freedom during the transient for other secondary tasks. Using simple Lyapunov arguments, it is possible to show that the law (2) is locally stable and that the secondary task does not impact the stability. If the secondary velocities $\bm{\sigma}$ are not compatible with the primary task, they are simply ignored and not executed, as the effect of the construction of the projector operator [41]. The main drawback of (2) is the presence of singularities for which a switching strategy is necessary around $\eta=0$:
$\bm{v}=\alpha(\eta)\,\bm{v}_{\eta}+\big{(}1-\alpha(\eta)\big{)}\,\bm{v}_{e}.$ (3)
For more details about the switching strategy, the interested reader could refer to the original paper [42].
@@ 4 Approach
Our objective is to exploit the control redundancy offered by the large projector formalism (2) to integrate DL and IL for direct VS. We exploit the great potential of state-of-the-art DL-based object detectors, which are treated as a backbone detector. Furthermore, our framework uses IL to look at previous demonstrations to overcome the limitations of DL-based detectors. The scheme further exploits IL to realize complex trajectories. In this section, we explain in detail the DL and IL learning components, and how they are combined using the large null-space projector control structure. More in detail, we first present and formulate the perception problem in Sec. 4.1, which regards the limitation of the backbone detectors in the control context; then, we present our solution to the problem introducing the imitation strategy (in Sec. 4.2) and the whole control scheme (Sec. 4.3) of our approach.
4.1 DL-based detection backbone and its limits
The DL-based backbone detector is a pre-trained NN that is fed with raw camera images and provides as output a measure of the object in the form of visual features
@@ 4.1 DL-based detection backbone and its limits
where $\bm{i}\in\mathbb{R}^{3wh}$ is a vectorized colored image with a size of $w\times h$ pixels, $\bm{m}$ the backbone model, and $\bm{\theta}\in\mathbb{R}^{p}$ denotes the parameters of the pre-trained model; the output $\bm{f}\in\mathbb{R}^{m}$ contains features of the detected objects, such as the corners of its bounding box measured on the image (see Fig. 2). Following the classic VS rationale, such features are compared with a set of desired values, denoted with $\bm{f}^{\ast}$, to provide a measure of the visual error, from which the control action evolves. Indeed, the desired set of features is obtained by the model fed with a reference desired image, i.e., $\bm{f}^{\ast}=\bm{m}_{\bm{\theta}}(\bm{i}^{\ast})$. Therefore, the visual error to be considered in the standard VS law (1) is computed as
@@
State-of-the-art systems, such as the YOLO algorithm [17] considered in our work, can perform high-frequency and robust feature detection. However, such features are not truly informative of the real pose of the observed object and, thus, not enough for full servoing of the camera pose. Typically, the bounding boxes detected by YOLO are rectangles centered on the image of the object of interest and aligned to the image borders. Such bounding boxes do not include any information about the object orientation, as shown in Fig. 2. An extended version of YOLO, YOLOv8 [43], computes bounding boxes that are oriented in the plane of the image. However, even in this case, the detected bounding box features do not contain information about the complete orientation of the detected object in all three rotation axes. Thus, they are insufficient for controlling the full 6D pose of the camera and its motion in the Cartesian space. One possible solution would be to refine the model $\bm{m}_{\bm{\theta}}$ by fine-tuning the parameters $\bm{\theta}$. However, this solution requires engineering work and the intervention of specialized scientists. Furthermore, the achievement of satisfactory results remains challenging. Our solution, instead, proposes to learn how to overcome these limitations of the DL backbone from given demonstrations of the desired task.
4.2 Overcoming the backbone limits through imitation
To overcome the limitations of DL-based backbone detectors like YOLO, we leverage the information contained in a set of human-demonstrated trajectories. In particular, the corrective action for servoing the 3D orientation can be imitated from demonstrations of the full VS behavior. Such demonstrations are contained in a dataset with this shape:
@@ 4.2 Overcoming the backbone limits through imitation
where the subscript $t$ and the superscript $n$ denote the $t$-th sample of the $n$-th demonstration, respectively; $N$ is the total number of demonstrations and $T$ is the length (expressed as the total number of samples) of each demonstration; $\bm{f}$ indicates the visual features vector (as defined in Sec. 4.1), and $\bm{p}\in\mathbb{R}^{3}$ is the robot’s end-effector’s position. The end-effector orientation $\bm{r}\in\mathbb{R}^{3}$ is obtained by projecting a unit quaternion $\bm{q}\in\mathbb{S}^{3}$ into the tangent space placed at the goal quaternion using the so-called logarithmic map [44]. Such quantities are obtained by showing the full desired behavior to the robot by using, e.g., teleoperation or kinesthetic teaching. During the collection of demonstrations, the object of interest is placed at a fixed location w.r.t. the robot, and the object is always maintained in the camera field of view.
Figure 3: The proposed framework for Direct Imitation Learning-based Visual Servoing (D-ILVS) exploits a backbone perception (a frozen DL network, implemented by YOLO) to extract features from raw images robustly and IL (implemented as a fine-tuned NODE network) to realize complex trajectories and overcome the limitation of the backbone perception. The large projection formulation merges the output of the backbone perception and imitation strategy in a closed-loop control law resulting in accurate and converging robot movements.
@@
The IL strategy is realized by augmenting the backbone model with additional layers. More in detail, we augment the perception backbone with a Neural Ordinary Differential Equation (NODE) solver [45] used for IL. The additional NODE layers are trained on the dataset ${\cal D}$. NODE has previously been used for IL [46] because it can be trained easily with only a few demonstrations, is extremely fast during inference, and exhibits accurate empirical performance for real-world full 6 degrees-of-freedom trajectory learning tasks [47]. The lack of mathematical stability guarantees of the trajectories predicted by a stand-alone NODE is addressed in our approach by the large projector formalism that guarantees that the position trajectory of the robot will not diverge, as discussed in Sec. 5. Hence, we do not use other alternatives such as [48] that assure stability but have a much slower inference speed [46].
@@
NODE assumes that the training data are instances of a nonlinear dynamical system that maps a generic input state $\bm{x}$ into an output that consists of its time derivative $\dot{\bm{x}}$. In our setting, the state at time sample $t$ is $\bm{x}_{t}=\left[\bm{f}_{t}^{\top},\bm{p}_{t}^{\top},\bm{r}_{t}^{\top}\right]^{\top}$. To accurately approximate the underlying dynamics, NODE optimizes a set of parameters $\bm{\vartheta}$ by minimizing a sum-of-square-error loss. It is worth mentioning that, having projected unit quaternions in the tangent space, we can readily use the dataset $\mathcal{D}$ as in (7) to train NODE. This is a common strategy in robotics [44, 49, 50, 51], which is also effective for NODE [46]. After training NODE, the rotation component is transformed back into unit quaternions using the so-called exponential map [44, 47]. While training NODE, in each iteration we extract from the $N$ demonstrations in $\mathcal{D}$ a short contiguous segment of length $T_{s}$, obtained by drawing from $\mathcal{D}$ elements at random temporal locations $T_{s}$, $T_{s}\ll T$ [46]. We then concatenate each element of $\mathcal{D}_{s}$ into the vectors $\bm{x}_{t}^{n}=\left[{\bm{f}_{t}^{n}}^{\top},{\bm{p}_{t}^{n}}^{\top},{\bm{r}_{t}^{n}}^{\top}\right]^{\top}$, $t=1,\ldots,T_{s}$, $n=1,\ldots,N$. Given the input vectors $\bm{x}_{t}^{n},\forall t,n$, NODE uses its internal neural network $\bm{n}_{\vartheta}$ (called target network) to produce derivatives of the input that are then numerically integrated to produce a predicted trajectory $\hat{\bm{x}}_{t}^{n},\forall t,n$. For training NODE, we used the mean squared error loss $\mathcal{L}$, defined as:
@@
that is initialized with the initial state of the system comprised of the output $\bm{f}_{0}$ of the pre-trained detection model (i.e., the YOLO backbone) and the initial robot’s position $\bm{p}_{0}$ and tangent space orientation $\bm{r}_{0}$. As output, it produces the robot velocity $\bm{\sigma}\in\mathbb{R}^{6}$, which is the time derivative of $\bm{p}$ and $\bm{r}$, imitating the complex trajectories demonstrated in the dataset. In subsequent steps of the robot’s motion, NODE evolves in an open-loop fashion its internal belief of the current state of $\hat{\bm{f}}$, $\hat{\bm{p}}$, and $\hat{\bm{r}}$, while keeping on predicting the velocity $\bm{\sigma}$.

### v2

# Imitation Learning-based Direct Visual Servoing using the Large Projection Formulation

@@ Abstract
Today robots must be safe, versatile, and user-friendly to operate in unstructured and human-populated environments. Dynamical system-based imitation learning enables robots to perform complex tasks stably and without explicit programming, greatly simplifying their real-world deployment. To exploit the full potential of these systems it is crucial to implement closed loops that use visual feedback. Vision permits to cope with environmental changes, but is complex to handle due to the high dimension of the image space. This study introduces a dynamical system-based imitation learning for direct visual servoing. It leverages off-the-shelf deep learning-based perception modules to extract robust features from the raw input image, and an imitation learning strategy to execute sophisticated robot motions. The learning blocks are integrated using the large projection task priority formulation. As demonstrated through extensive experimental analysis, the proposed method realizes complex tasks with a robotic manipulator.
@@
Figure 1: Our work combines off-the-shelf deep learning strategies to detect objects in the clutter, and imitation learning to realize complicated trajectories, e.g., dropping a cube into a cup on an untidy table. The large projection formulation combines the two machine learning components and ensures convergence to a given target.
@@ 1 Introduction
Modern robots must be accessible to everyone, as they are rapidly spreading in everyday life environments, from industries [1] to hotels [2] and hospitals [3]. It is expected that a growing number of inexperienced end-users, like children, patients, or elderly people, ask for robots with easy-to-use, friendly, and modular interfaces, endowed with adaptive skills. To meet these requirements, recent advancements in robotics demonstrate the great potential of Machine Learning ( Machine Learning (ML)).
@@ 2 Related work
One possibility consists of coupling perception and control together in end-to-end learning fashion [29, 30, 31]. However, end-to-end methods cannot guarantee stability properties and robustness to disturbances. This is particularly challenging if the robot needs to operate in dynamically changing environments and/or close to the human. Coarse-to-fine imitation learning [32] combines closed- and open-loop execution to perform complex manipulation tasks. One way to ensure stability is to maintain the formal structure of the visual controller, e.g., preserving the VS formalism. In the context of our work, Direct (DVS) is particularly interesting because it implements VS using direct image measurement, avoiding explicit feature extraction. Examples are VS schemes that use photometric moments [33], pixel luminance [34], histograms [35], or Gaussian mixtures [36] as control feedback. However, in these approaches, the potential of DL is not fully exploited. In [37], an artificial Neural Network (NN) is trained using the knowledge of VS to produce geometrically interpretable features. In [38], an autoencoder is learned to reduce the dimensionality of the image space, and an interaction matrix is directly computed from the network using auto differentiation and utilized in a VS law. Liu et al. [39] simplify the YOLO architecture to speed up object detection, while Luo et al. [40] propose a top-down feature detection network, and both use the predicted features in a VS scheme. These approaches perform well with VS tasks from images without performing classical feature extraction but do not execute complex movements. Our work, instead, provides a unified solution that generates sophisticated robot motion in addition to enhancing the robustness of the perception module.
This paper presents an approach to solve the perception problem and the generation of complex trajectories simultaneously, in the context of vision-based controllers. The proposed framework overcomes the limitations of the literature by proposing an Imitation Learning-based Direct Visual Servoing (IL-DVS) strategy that integrates off-the-shelf DL-based perception modules with IL in a control theoretic framework.
@@ 3 Background
Image-based VS [10] is a more than twenty-year-old established technique to regulate a camera to a desired pose through visual information. The most basic law computes 6D velocity commands as $\bm{v}=-\lambda\hat{\bm{L}}^{+}\bm{e}$ [10], zeroing an error $\bm{e}\in\mathbb{R}^{f}$ defined on the image; $\lambda$ is a scalar gain; $\hat{\bm{L}}^{+}\in\mathbb{R}^{f\times 6}$ is the pseudo-inverse of the so-called interaction matrix, relating the camera velocity to the time derivative of the visual feedback. The interaction matrix normally relies on the camera parameters that can be obtained through calibration procedures, and other information like the features’ depth that needs to be estimated or approximated. Indeed, the hat over the matrix indicates the approximation due to unknown $3$D parameters.
(a) Side view.
(b) Oblique view.
(c) Top view.
Figure 2: State-of-the-art DL-based systems like YOLO can be used to detect the features of an object of interest on a raw monocular image robustly. Examples of features are the vertices (denoted with ‘C1’, ‘C2’, ‘C3’, and ‘C4’) of the bounding box detected around the image of a cup. However, such detection systems fail to capture the correct object orientation. In the three snapshots, YOLO provides very similar feature values that correspond to three very different relative camera-object orientations producing a side (a), oblique (b), and top view (c) of the cup.
@@
$\bm{v}_{e}=-\lambda\hat{\bm{L}}^{+}\bm{e}+\bm{P}\bm{\sigma}.$ (1)
The matrix $\bm{P}=\bm{I}_{6}-\hat{\bm{L}}^{+}\hat{\bm{L}}$ is a null-space projector, and $\bm{\sigma}\in\mathbb{R}^{6}$ is the desired velocity realizing the secondary task. The main limitation of (1) is that, in normal working conditions, the dimension of the feedback is always greater or equal to the dimension of the task (i.e., $f\geq 6$[10]. Under these circumstances, there is not much room in the null space of the primary task to execute any other secondary task. Therefore, it has been proposed to use the norm of the error $\eta=\|\bm{e}\|$ in the primary task [42] and to realize a prioritized control scheme in the following form
@@
where, similarly to the classic VS scheme (1), $\bm{P}_{\eta}$ is a null-space projector and $\bm{\sigma}\in\mathbb{R}^{6}$ is the desired velocity realizing the secondary task. The matrices of interest can be retrieved in closed form [42], and are reported here for convenience:
$\hat{\bm{L}}_{\eta}^{+}=\eta\frac{\bm{\hat{L}}^{\top}\bm{e}}{\bm{e}^{\top}\bm{\hat{L}}\bm{\hat{L}}^{\top}\bm{e}}\quad\text{and}\quad\bm{P}_{\eta}=\bm{I}_{6}-\frac{\bm{\hat{L}}^{\top}\bm{e}\,\bm{e}^{\top}\bm{\hat{L}}}{\bm{e}^{\top}\bm{\hat{L}}\bm{\hat{L}}^{\top}\bm{e}}.$
Note that the control law (2) drives the system towards the original desired behavior, as $\bm{e}\to\bm{0}$ for $\eta\to 0$. The operator $\bm{P}_{\eta}$ is called large projector as the law (2) ensures enough room in the null space to achieve secondary tasks. In practice, regulating the scalar norm instead of the vector error releases degrees of freedom during the transient for other secondary tasks. Using simple Lyapunov arguments, it is possible to show that the law (2) is locally stable and that the secondary task does not impact the stability. If the secondary velocities $\bm{\sigma}$ are not compatible with the primary task, they are simply ignored and not executed, as the effect of the construction of the projector operator [41]. In (2) the presence of singularities requires the use of a switching strategy around $\eta=0$:
$\bm{v}=\alpha(\eta)\,\bm{v}_{\eta}+\big{(}1-\alpha(\eta)\big{)}\,\bm{v}_{e}$ (3)
where $\alpha$ is a scalar variable smoothly changing from $1$ to 00 in the vicinity of $\eta=0$, allowing a switch to the classical law (1) from (2[42].
@@ 4 Approach
Our objective is to exploit the control redundancy offered by the large projector formalism (2) to integrate DL and IL for direct VS. We exploit the great potential of state-of-the-art DL-based object detectors, which are treated as an underlying raw detector. Furthermore, our framework uses IL to look at previous demonstrations to overcome the limitations of DL-based detectors. The scheme further exploits IL to realize complex trajectories. In this section, we explain in detail the DL and IL learning components, and how they are combined using the large null-space projector control structure. More in detail, we first present and formulate the perception problem in Sec. 4.1, which regards the limitation of the used detectors in the control context; then, we present our solution to the problem introducing the imitation strategy (in Sec. 4.2) and the whole control scheme (Sec. 4.3) of our approach.
4.1 DL-based detection and its limits
The DL-based detector is a pre-trained NN that is fed with raw camera images and provides as output a measure of the object in the form of visual features
@@ 4.1 DL-based detection and its limits
where $\bm{i}\in\mathbb{R}^{3wh}$ is a vectorized colored image with a size of $w\times h$ pixels, $\bm{m}$ is the detector model, and $\bm{\theta}\in\mathbb{R}^{p}$ denotes the parameters of the pre-trained model; the output $\bm{f}\in\mathbb{R}^{m}$ contains features of the detected objects, such as the corners of its bounding box measured on the image (see Fig. 2). Following the classic VS rationale, such features are compared with a set of desired values, denoted with $\bm{f}^{\ast}$, to provide a measure of the visual error, from which the control action evolves. Indeed, the desired set of features is obtained by the model fed with a reference desired image $\bm{i}^{\ast}$, i.e., $\bm{f}^{\ast}=\bm{m}_{\bm{\theta}}(\bm{i}^{\ast})$. Therefore, the visual error to be considered in the standard VS law (1) is computed as
@@
State-of-the-art systems, such as the YOLO algorithm [17] considered in our work, can perform high-frequency and robust feature detection. However, such features are not truly informative of the real pose of the observed object and, thus, not enough for full servoing of the camera pose. Typically, the bounding boxes detected by YOLO are rectangles centered on the image of the object of interest and aligned to the image borders. Such bounding boxes do not include any information about the object orientation, as shown in Fig. 2. An extended version of YOLO, YOLOv8 [43], computes bounding boxes that are oriented in the plane of the image. However, even in this case, the detected bounding box features do not contain information about the complete orientation of the detected object in all three rotation axes. Thus, they are insufficient for controlling the full 6D pose of the camera and its motion in the Cartesian space. One possible solution would be to refine the model $\bm{m}_{\bm{\theta}}$ by fine-tuning the parameters $\bm{\theta}$. However, this solution requires engineering work and the intervention of specialized scientists. Furthermore, the achievement of satisfactory results remains challenging. Our solution, instead, proposes to learn how to overcome these limitations of the off-the-shelf DL detection module from given demonstrations of the desired task.
Figure 3: The proposed framework for IL-DVS exploits a detection model (a frozen DL network, implemented by YOLO) to extract features from raw images robustly and IL (implemented as a fine-tuned NODE network) to realize complex trajectories and overcome the limitation of the detection model. The large projection formulation merges the output of the detection and imitation strategy in a closed-loop control law resulting in accurate and converging robot movements.
4.2 Overcoming the detection limits through imitation
To overcome the limitations of DL-based detectors like YOLO, we leverage the information contained in a set of human-demonstrated trajectories. In particular, the corrective action for servoing the 3D orientation can be imitated from demonstrations of the full VS behavior. Such demonstrations are contained in a dataset with this shape:
@@ 4.2 Overcoming the detection limits through imitation
where the subscripts $t$ and $n$ denote the $t$-th sample of the $n$-th demonstration, respectively; $N$ is the total number of demonstrations and $T$ is the length (expressed as the total number of samples) of each demonstration; $\bm{f}$ indicates the visual features vector (as defined in Sec. 4.1), and $\bm{p}\in\mathbb{R}^{3}$ is the robot’s end-effector’s position. The end-effector orientation $\bm{r}\in\mathbb{R}^{3}$ is obtained by projecting a unit quaternion $\bm{q}\in\mathbb{S}^{3}$ into the tangent space placed at the goal quaternion using the so-called logarithmic map [44]. Such quantities are obtained by showing the full desired behavior to the robot by using, e.g., teleoperation or kinesthetic teaching. During the collection of demonstrations, the object of interest is placed at a fixed location w.r.t. the robot, and the object is always maintained in the camera field of view.
@@
The IL strategy is realized by augmenting the detection model with additional layers. More in detail, we augment the detector architecture with a Neural Ordinary Differential Equation (NODE) solver [45] used for IL. The additional NODE layers are trained on the dataset ${\cal D}$. NODE has previously been used for IL [46] because it can be trained easily with only a few demonstrations, is extremely fast during inference, and exhibits accurate empirical performance for real-world full 6 degrees-of-freedom trajectory learning tasks [47]. The lack of mathematical stability guarantees of the trajectories predicted by a stand-alone NODE is addressed in our approach by the large projector formalism that guarantees that the position trajectory of the robot will not diverge, as discussed in Sec. 5. Hence, we do not use other alternatives such as [48] that assure stability but have a much slower inference speed [46].
@@
NODE assumes that the training data are instances of a nonlinear dynamical system that maps a generic input state $\bm{x}$ into an output that consists of its time derivative $\dot{\bm{x}}$. In our setting, the state at time sample $t$ is $\bm{x}_{t}=\left(\bm{f}_{t}^{\top},\bm{p}_{t}^{\top},\bm{r}_{t}^{\top}\right)^{\top}$. To accurately approximate the underlying dynamics, NODE optimizes a set of parameters $\bm{\vartheta}$ by minimizing a sum-of-square-error loss. It is worth mentioning that, having projected unit quaternions in the tangent space, we can readily use the dataset $\mathcal{D}$ as in (7) to train NODE. This is a common strategy in robotics [44, 49, 50, 51], which is also effective for NODE [46]. After training NODE, the rotation component is transformed back into unit quaternions using the so-called exponential map [44, 47]. While training NODE, in each iteration we extract from the $N$ demonstrations in $\mathcal{D}$ a short contiguous segment of length $T_{s}$, obtained by drawing from $\mathcal{D}$ elements at random temporal locations $T_{s}$, $T_{s}\ll T$ [46]. We then concatenate each element of $\mathcal{D}_{s}$ into the vectors $\bm{x}_{t}^{n}=\left({\bm{f}_{t}^{n}}^{\top},{\bm{p}_{t}^{n}}^{\top},{\bm{r}_{t}^{n}}^{\top}\right)^{\top}$, $t=1,\ldots,T_{s}$, $n=1,\ldots,N$. Given the input vectors $\bm{x}_{t}^{n},\forall t,n$, NODE uses its internal neural network $\bm{n}_{\vartheta}$ (called target network) to produce derivatives of the input that are then numerically integrated to produce a predicted trajectory $\hat{\bm{x}}_{t}^{n},\forall t,n$. For training NODE, we used the mean squared error loss $\mathcal{L}$, defined as:
@@
that is initialized with the initial state of the system comprised of the output $\bm{f}_{0}$ of the pre-trained detection model (i.e., YOLO) and the initial robot’s position $\bm{p}_{0}$ and tangent space orientation $\bm{r}_{0}$. As output, it produces the robot velocity $\bm{\sigma}\in\mathbb{R}^{6}$, which is the time derivative of $\bm{p}$ and $\bm{r}$, imitating the complex trajectories demonstrated in the dataset. In subsequent steps of the robot’s motion, NODE evolves in an open-loop fashion its internal belief of the current state of $\hat{\bm{f}}$, $\hat{\bm{p}}$, and $\hat{\bm{r}}$, while keeping on predicting the velocity $\bm{\sigma}$.